# On algebraic properties of some non-prime ideals of collections of cells

### v2

@@ 1.3. The coordinate ring attached to a collection of cells
For utility, we state the following result about the Cohen-Macaulay property of the simple collection of cells, that in literature we found stated in this form only in the case $\mathcal{P}$ is polyomino (see [18, Theorem 2.1] and [19, Corollary 3.3]). For the sake of completeness, we also provide its proof.
@@
By [5, Theorem 3.3] we know that $I_{\mathcal{P}}=J_{\mathcal{P}}$ where $J_{\mathcal{P}}$ is the toric ring of a bipartite weakly chordal graph. In particular $I_{\mathcal{P}}$ is prime and, by [27], $I_{\mathcal{P}}$ has a squarefree quadratic Gröbner basis. A toric ring whose toric ideal admits a squarefree initial ideal is normal, see [16, Corollary 4.26] and by a theorem of Hochster (see [2, Theorem 6.3.5]), a normal toric ring is Cohen–Macaulay. Finally, by [28, Theorem 3.5] and [28, Corollary 3.6] we know that $I_{\mathcal{P}}=I_{\Lambda}$ for a saturated lattice $\Lambda$ such that $\operatorname{rank}_{\mathbb{Z}}(\Lambda)=|\mathcal{P}|$. Hence $\operatorname{height}(I_{\mathcal{P}})=|\mathcal{P}|$ (see [16, Problem 3.12]) and so $\dim K[\mathcal{P}]=|V(\mathcal{P})|-|\mathcal{P}|$ (see also [37, Corollary 3.1.7]). ∎
@@
Figure 4. A closed path with two $L$-configurations and a ladder of four steps.
@@
A path of five cells $C_{1},C_{2},C_{3},C_{4},C_{5}$ of $\mathcal{P}$ is called an L-configuration if the two sequences $C_{1},C_{2},C_{3}$ and $C_{3},C_{4},C_{5}$ go in two orthogonal directions. A set =$\mathcal{B}=\{\mathcal{B}_{i}\}_{i=1,\dots,n}$ of maximal horizontal (or vertical) blocks of rank at least two, with $V(\mathcal{B}_{i})\cap V(\mathcal{B}_{i+1})=\{a_{i},b_{i}\}$ and $a_{i}\neq b_{i}$ for all $i=1,\dots,n-1$, is called a ladder of $n$ steps if $[a_{i},b_{i}]$ is not on the same edge interval of $[a_{i+1},b_{i+1}]$ for all $i=1,\dots,n-2$. For instance, in Figure 4 there is a closed path having two $L$-configurations and a ladder of four steps. It is proved in [3, Section 6] that a closed path has no zig-zag walks if and only if it contains an $L$-configuration or a ladder of at least three steps. Furthermore, in [3, Theorem 6.2] the authors characterize the primality of a closed path $\mathcal{P}$ proving that $I_{\mathcal{P}}$ is prime if and only if $\mathcal{P}$ does not contain any zig-zag walk.
1.4. Some results in Commutative Algebra.
In this subsection we recall some well-known results in Commutative Algebra that we use extensively throughout the paper. We recall, for completeness, that $\operatorname{depth}(M)\leq\dim(M)$ holds for every $R$-module $M$ over a Noetherian ring $R$, and $M$ is Cohen-Macaulay when equality holds. About some notations we use throughout the paper, if $S$ is a polynomial ring, $\prec$ is a monomial order on it and $f\in S$ is a polynomial, we denote by $\operatorname{in}_{\prec}(f)$, or $\operatorname{in}(f)$ if the order is clear in the context, the initial monomial of $f$ with respect to $\prec$. As usual, if $S=K[x_{1},\ldots,x_{n}]$ and $m$ is a monomial, then $\operatorname{supp}(m)=\{x_{i}\in S\mid x_{i}\text{ divides }m\}$.
We start by providing some basics on the Hilbert function and the Hilbert-Poincaré series of a graded $K$-algebra $R/I$. Let $R$ be a graded $K$-algebra and $I$ be an homogeneous ideal of $R$. Then $R/I$ has a natural structure of graded $K$-algebra as $\bigoplus_{k\in\mathbb{N}}(R/I)_{k}$. The numerical function $\mathrm{H}_{R/I}:\mathbb{N}\to\mathbb{N}$ with $\mathrm{H}_{R/I}(k)=\dim_{K}(R/I)_{k}$ is called the Hilbert function of $R/I$. The formal power series $\mathrm{HP}_{R/I}(t)=\sum_{k\in\mathbb{N}}\mathrm{H}_{R/I}(k)t^{k}$ is called the Hilbert-Poincaré series of $R/I$. It is known by Hilbert-Serre theorem that there exists a polynomial $h(t)\in\mathbb{Z}[t]$ with $h(1)\neq 0$ such that $\mathrm{HP}_{R/I}(t)=\frac{h(t)}{(1-t)^{d}}$, where $d$ is the Krull dimension of $R/I$. The polynomial $h(t)$ is called h-polynomial of $R/I$. Moreover, if $R/I$ is Cohen-Macaulay then $\mathrm{reg}(R/I)=\deg h(t)$. Recall also that if $S=K[x_{1},\ldots,x_{n}]$ then $\mathrm{HP}_{S}(t)=\frac{1}{(1-t)^{n}}$. In some parts of the paper, if $K[\mathcal{P}]$ is the coordinate ring of a collection of cells, we use the notation $h_{K[\mathcal{P}]}$ for the $h$-polyonimial of the Hilbert-Poincaré series of $K[\mathcal{P}]$.
Proposition 1.5.
If $I$ is a homogeneous ideal of a graded ring $R$ then:
  1. (1)
    $R/I$ and $R/\mathrm{in}(I)$ have the same Hilbert function for all monomial orders (see [37, Corollary 3.3.15]);
  2. (2)
    If $R/\mathrm{in}(I)$ is Cohen–Macaulay (resp. Gorenstein), then $R/I$ is Cohen–Macaulay (resp. Gorenstein) (see [37, Proposition 9.6.17]).
From [9] we deduce that the converse for the Cohen-Macaulayness in (2) of Proposition 1.5 holds for a homogeneous ideal whose initial ideal is squarefree.
Proposition 1.6 ([9]).
Let $I$ be a homogeneous ideal in a polynomial ring $S$ and suppose $\mathrm{in}(I)$ is radical with respect to some term order. Then $\mathrm{reg}(S/I)=\mathrm{reg}(S/\mathrm{in}(I))$ and $\mathrm{depth}(S/I)=\mathrm{depth}(S/\mathrm{in}(I))$. Furthermore, $S/I$ is Cohen-Macaulay if and only if $S/\mathrm{in}(I)$ is Cohen-Macaulay.
For the following result, see Lemma 5.1.2 and Lemma 5.1.3 of [37].
Proposition 1.7.
Let $R$ be a graded $K$-algebra and $I$ be a graded ideal of $R$. Let $q$ be an homogeneous element of $R$ of degree $m$ and let
$0\longrightarrow R/(I:q)\longrightarrow R/I\longrightarrow R/(I,q)\longrightarrow 0$
be a short exact sequence. Then $\mathrm{HP}_{R/I}(t)=\mathrm{HP}_{R/(I,q)}(t)+t^{m}\mathrm{HP}_{R/(I:q)}(t)$.
For the following result, see [37, Lemma 5.1.11].
Proposition 1.8.
Let $A$ and $B$ be standard graded $K$-algebras over a field $K$. Then $\mathrm{HP}_{A\otimes_{K}B}(t)=\mathrm{HP}_{A}(t)\cdot\mathrm{HP}_{B}(t)$. In particular $\mathrm{dim}(A\otimes_{K}B)=\mathrm{dim}(A)+\mathrm{dim}(B)$.
For the following result, see Lemma 3.1.34, Lemma 3.1.35 and Proposition 11.5.8 of [37].
Proposition 1.9.
Let $A$ and $B$ be standard graded $K$-algebras over a field $K$. Then $\mathrm{depth}(A\otimes_{K}B)=\mathrm{depth}(A)+\mathrm{depth}(B)$. In particular $A\otimes_{K}B$ is Cohen-Macaulay (Gorenstein) if and only if $A$ and $B$ are Cohen-Macaulay (Gorenstein).
For the following result, see [2, Proposition 1.2.9].
Proposition 1.10.
Let $0\longrightarrow N_{1}\longrightarrow N_{2}\longrightarrow N_{3}\longrightarrow 0$ be a short exact sequence of finitely generated $R$-module over a Noetherian ring $R$. Then:
  • $\mathrm{depth}(N_{2})\geq\min\{\mathrm{depth}(N_{3}),\mathrm{depth}(N_{1})\}$.
  • $\mathrm{depth}(N_{1})\geq\min\{\mathrm{depth}(N_{2}),\mathrm{depth}(N_{3})+1\}$.
  • $\mathrm{depth}(N_{3})\geq\min\{\mathrm{depth}(N_{1})-1,\mathrm{depth}(N_{2})\}$.
@@ 2. Hilbert-Poincaré series of simple collections of cells
In this section we present some results on simple collections of cells. From [6, Theorem 3.3] and [27] we know that, if $\mathcal{P}$ is a simple collection of cells, the set of generators of $I_{\mathcal{P}}$ forms the reduced Gröbner basis of $I_{\mathcal{P}}$ with respect to a suitable monomial order. The monomial order provided in [27] is fundamental for our purposes because it allows us to write $S_{\mathcal{P}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}})$ as $K$-tensor product of suitable $K$-algebras coming from the connected components. We start proving the following Lemma.
@@
Let $\mathcal{P}$ be a simple collection of cells. According to the references [5, Theorem 3.3] and [27], $I_{\mathcal{P}}$ possesses a quadratic Gröbner basis with respect to a certain monomial order. We fix such a monomial order, then for any pair of inner 2-minors $f$ and $g$, their $S$-polynomial $S(f,g)$ reduces to 0. Suppose that $f$ and $g$ are inner 2-minors of $\mathcal{P}_{i}$ and $\mathcal{P}_{j}$ respectively, where $i\neq j$. Now we have two cases: when $f$ and $g$ are binomials associated with inner intervals that are disjoint or they share a vertex. For the first, we have $\gcd(\mathop{\rm in}\nolimits(f),\mathop{\rm in}\nolimits(g)=1$ trivially. For the second case, since the inner 2-minors $f$ and $g$ are of the form as shown in Figure 5 and $S(f,g)$ reduces to 00, so by Lemma 2.1 we obtain $\gcd(\mathop{\rm in}\nolimits(f),\mathop{\rm in}\nolimits(g))=1$ as well. So, for all $i,j\in[n]$ with $i\neq j$ and for all $f,g$ inner 2-minors of $I_{\mathcal{P}}$ such that $f$ is related to an inner interval of $\mathcal{P}_{i}$ and $g$ is related to an inner interval of $\mathcal{P}_{j}$, we have $\gcd(\mathop{\rm in}\nolimits(f),\mathop{\rm in}\nolimits(g))=1$. As a consequence, if $v\in V(\mathcal{P}_{i})\cap V(\mathcal{P}_{j})$ for some $i,j\in[n]$ with $i\neq j$, and $x_{v}\in\mathrm{supp}(\mathop{\rm in}\nolimits(f))$, for some inner 2-minor $f$ related to an inner interval of $\mathcal{P}_{i}$, then $x_{v}\notin\mathrm{supp}(\mathop{\rm in}\nolimits(g))$, for all inner 2-minors $g$ related to the inner intervals of $\mathcal{P}_{j}$. Let $A_{\mathcal{P}}=\{v\in V(\mathcal{P})\mid v\in V(\mathcal{P}_{i})\cap V(\mathcal{P}_{j}),\text{ for some }i,j\in[n],i\neq j\}$. Let $C_{\mathcal{P}}$ be the set of vertices $v\in A_{\mathcal{P}}$ such that $x_{v}\notin\mathrm{supp}(\mathop{\rm in}\nolimits(f))$, for all inner 2-minor $f$ related to an inner interval of $\mathcal{P}$. For each $i\in[n]$, define $D_{\mathcal{P}_{i}}$ the set of vertices $v\in V(\mathcal{P}_{i})\cap A_{\mathcal{P}}$ such that $x_{v}\notin\mathrm{supp}(\mathop{\rm in}\nolimits(f)),\text{ for all inner 2-minor }f\text{ related to an inner interval of }\mathcal{P}_{i}$. Denote $d_{i}=|D_{\mathcal{P}_{i}}|$. Set, for all $i\in[n]$, $S^{\prime}_{\mathcal{P}_{i}}=K[x_{a}:a\in V(\mathcal{P}_{i})\setminus D_{\mathcal{P}_{i}}]$. Then we have:
@@
Observe that for all $i\in[n]$ we have $S_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})=S^{\prime}_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})\otimes_{K}K[x_{v}\mid v\in D_{\mathcal{P}_{i}}]$. In particular, by Proposition 1.8, $\mathrm{HP}_{S^{\prime}_{\mathcal{P}_{i}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}_{i}})}(t)=(1-t)^{d_{i}}\mathrm{HP}_{K[\mathcal{P}_{i}]}(t)$, and since $\mathcal{P}_{i}$ is a simple polyomino, by Proposition 1.3 we have $\mathrm{HP}_{S^{\prime}_{\mathcal{P}_{i}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}_{i}})}(t)=\frac{h_{\mathcal{P}_{i}}(t)}{(1-t)^{|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-d_{i}}}$. Note that $\sum_{i=1}^{n}|V(\mathcal{P}_{i})|=|V(\mathcal{P})|+|A_{\mathcal{P}}|$ and $\sum_{i=1}^{n}|\mathcal{P}_{i}|=|\mathcal{P}|$. Therefore, by Proposition 1.8, we have:
@@
By the previous result, we argue that, if $\mathcal{P}$ is a simple collection of cells, the study of the $h$-polynomial of $K[\mathcal{P}]$ in terms of the switching rook polynomial of $\mathcal{P}$ can be reduced to investigate the Hilbert-Poincaré series of its connected components. In particular, if for all simple polyominoes [29, Conjecture 3.2] is true, then the same holds for all simple collections of cells.
@@
Let $\mathcal{P}$ be a simple collection of cells where every connected component $\mathcal{P}_{i}$ is a parallelogram polyomino. Then, by [29, Theorem 3.5], $h_{K[\mathcal{P}_{i}]}(t)$ is the switching rook polynomial of $\mathcal{P}_{i}$ for all $i\in[n]$, so $h_{K[\mathcal{P}]}$ is the switching rook polynomial of $\mathcal{P}$ by Corollary 2.3.
As a consequence, we generalize [31, Theorem 3.13 and Theorem 4.2] to simple thin collections of cells (see Figure 6 for an example). This is useful in the next sections.
@@
Figure 6. A simple thin polyomino and a simple thin collection of cells.
@@
As a consequence of Corollary 2.3 and [31, Theorem 3.13], we get that $h_{K[\mathcal{P}]}$ is the rook polynomial of $\mathcal{P}$. Moreover, $\operatorname{reg}(K[\mathcal{P}])=r(\mathcal{P})$ since $K[\mathcal{P}]$ is a Cohen-Macaulay ring. Now we prove the result about the Gorensteiness. Let $h_{K[\mathcal{P}]}(t)=r_{\mathcal{P}}(t)=\sum_{k=0}^{s}r_{k}t^{k}$, where $s=r(\mathcal{P})$. Since $K[\mathcal{P}]$ is a Cohen-Macaulay domain, it is known from [35] that $K[\mathcal{P}]$ is Gorenstein if and only if $r_{i}=r_{s-i}$ for all $i=0,\dots,s$, that is $h_{K[\mathcal{P}]}(t)$ is palindromic. Assume that $\mathcal{P}$ consists of $n$ connected components and denote by $h_{K[\mathcal{P}_{i}]}(t)$ the $h$-polynomial of $K[\mathcal{P}_{i}]$ for all $i\in[n]$. Suppose that $\mathcal{P}$ has the $S$-property and we prove that $K[\mathcal{P}]$ is Gorenstein. Since $\mathcal{P}$ has the $S$-property, then each connected component $\mathcal{P}_{i}$ has the $S$-property as well. By [31, Theorem 4.2], $h_{K[\mathcal{P}_{i}]}(t)$ is palindromic for all $i\in[n]$, so $h_{K[\mathcal{P}]}(t)$ is palindromic as well, since $h_{K[\mathcal{P}]}(t)=\prod_{i=1}^{n}h_{K[\mathcal{P}_{i}]}(t)$ and the product of palindromic polynomials is palindromic. Hence it follows by [35] that $K[\mathcal{P}]$ is Gorenstein. Now, suppose that $K[\mathcal{P}]$ is Gorenstein and we show that $\mathcal{P}$ has the $S$-property. In particular the rook polynomial of $\mathcal{P}$, that is $h_{K[\mathcal{P}]}$, is palindromic and observe also that $r_{0}=1$ and $r_{1}=|\mathcal{P}|$. We assume by contradiction that $\mathcal{P}$ does not have the $S$-property. For all $i\in[n]$, let $h_{K[\mathcal{P}_{i}]}(t)=\sum_{j=0}^{s_{i}}r_{j}^{(i)}t^{j}$ be the $h$-polynomial of $K[\mathcal{P}_{i}]$, where $s_{i}=r(\mathcal{P}_{i})$ and observe that $r_{1}^{(i)}=|\mathcal{P}_{i}|$. Looking at the proof of case (b)$\Rightarrow$(c) of [31, Theorem 4.2], for all $k\in[n]$ such that $\mathcal{P}_{k}$ has not the $S$-property we have $r_{s_{k}}>1$ or $r_{s_{k}-1}>|\mathcal{P}_{k}|$. Therefore, it follows easily that $r_{s}>1$ or $r_{s-1}>\sum_{i=1}^{n}|\mathcal{P}_{i}|=|\mathcal{P}|$, that is a contradiction, since $h_{K[\mathcal{P}]}(t)$ is palindromic. ∎

### v3

@@ 1.3. The coordinate ring attached to a collection of cells
For utility, we state the following result about the Cohen-Macaulay property of the simple collection of cells, that in the literature we found stated in this form only in the case $\mathcal{P}$ is polyomino (see [19, Theorem 2.1] and [20, Corollary 3.3]). For the sake of completeness, we also provide its proof.
@@
By [6, Theorem 3.3] we know that $I_{\mathcal{P}}=J_{\mathcal{P}}$ where $J_{\mathcal{P}}$ is the toric ring of a weakly chordal bipartite graph. In particular $I_{\mathcal{P}}$ is prime and, by [28], $I_{\mathcal{P}}$ has a squarefree quadratic Gröbner basis. A toric ring whose toric ideal admits a squarefree initial ideal is normal, see [17, Corollary 4.26] and by a theorem of Hochster (see [3, Theorem 6.3.5]), a normal toric ring is Cohen–Macaulay. Finally, by [29, Theorem 3.5] and [29, Corollary 3.6] we know that $I_{\mathcal{P}}=I_{\Lambda}$ for a saturated lattice $\Lambda$ such that $\operatorname{rank}_{\mathbb{Z}}(\Lambda)=|\mathcal{P}|$. Hence $\operatorname{height}(I_{\mathcal{P}})=|\mathcal{P}|$ (see [17, Problem 3.12]) and so $\dim K[\mathcal{P}]=|V(\mathcal{P})|-|\mathcal{P}|$ (see also [37, Corollary 3.1.7]). ∎
@@
Figure 3. Examples of closed paths.
@@
A path of five cells $C_{1},C_{2},C_{3},C_{4},C_{5}$ of $\mathcal{P}$ is called an L-configuration if the two sequences $C_{1},C_{2},C_{3}$ and $C_{3},C_{4},C_{5}$ go in two orthogonal directions. A set =$\mathcal{B}=\{\mathcal{B}_{i}\}_{i=1,\dots,n}$ of maximal horizontal (or vertical) blocks of rank at least two, with $V(\mathcal{B}_{i})\cap V(\mathcal{B}_{i+1})=\{a_{i},b_{i}\}$ and $a_{i}\neq b_{i}$ for all $i=1,\dots,n-1$, is called a ladder of $n$ steps if $[a_{i},b_{i}]$ is not on the same edge interval of $[a_{i+1},b_{i+1}]$ for all $i=1,\dots,n-2$. For instance, in Figure 3a there is a closed path having two $L$-configurations and a ladder of four steps. It is proved in [4, Section 6] that a closed path has no zig-zag walks if and only if it contains an $L$-configuration or a ladder of at least three steps. Furthermore, in [4, Theorem 6.2] the authors characterize the primality of a closed path $\mathcal{P}$ proving that $I_{\mathcal{P}}$ is prime if and only if $\mathcal{P}$ does not contain any zig-zag walk.
@@ 2. Hilbert-Poincaré series of simple collections of cells
In this section we present some results on simple collections of cells. From [7, Theorem 3.3] and [28] we know that, if $\mathcal{P}$ is a simple collection of cells, the set of generators of $I_{\mathcal{P}}$ forms the reduced Gröbner basis of $I_{\mathcal{P}}$ with respect to a suitable monomial order. The monomial order provided in [28] is fundamental for our purposes because it allows us to write $S_{\mathcal{P}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}})$ as $K$-tensor product of suitable $K$-algebras coming from the connected components.
About some notations we use throughout the paper, if $S$ is a polynomial ring, $\prec$ is a monomial order on it and $f\in S$ is a polynomial, we denote by $\operatorname{in}_{\prec}(f)$, or $\operatorname{in}(f)$ if the order is clear in the context, the initial monomial of $f$ with respect to $\prec$. As usual, if $S=K[x_{1},\ldots,x_{n}]$ and $m$ is a monomial, then $\operatorname{supp}(m)=\{x_{i}\in S\mid x_{i}\text{ divides }m\}$. We start proving the following Lemma.
@@
Let $R$ be a graded $K$-algebra and $I$ be a homogeneous ideal of $R$. The formal power series $\mathrm{HP}_{R/I}(t)=\sum_{k\in\mathbb{N}}\dim_{K}(R/I)_{k}t^{k}$ is called the Hilbert-Poincaré series of $R/I$. It is known by Hilbert-Serre theorem that there exists a polynomial $h(t)\in\mathbb{Z}[t]$ with $h(1)\neq 0$ such that $\mathrm{HP}_{R/I}(t)=\frac{h(t)}{(1-t)^{d}}$, where $d$ is the Krull dimension of $R/I$. The polynomial $h(t)$ is called h-polynomial of $R/I$. Moreover, if $R/I$ is Cohen-Macaulay then $\mathrm{reg}(R/I)=\deg h(t)$. Recall also that if $S=K[x_{1},\ldots,x_{n}]$ then $\mathrm{HP}_{S}(t)=\frac{1}{(1-t)^{n}}$. In some parts of the paper, if $K[\mathcal{P}]$ is the coordinate ring of a collection of cells, we use the notation $h_{K[\mathcal{P}]}$ for the $h$-polyonimial of the Hilbert-Poincaré series of $K[\mathcal{P}]$. Once we set the previous notation, let us prove the main result of this section.
@@
Let $\mathcal{P}$ be a simple collection of cells. According to references [6, Theorem 3.3] and [28], $I_{\mathcal{P}}$ possesses a quadratic Gröbner basis with respect to a certain monomial order. We fix such a monomial order, then for any pair of inner 2-minors $f$ and $g$, their $S$-polynomial $S(f,g)$ reduces to 0. Suppose that $f$ and $g$ are inner 2-minors of $\mathcal{P}_{i}$ and $\mathcal{P}_{j}$ respectively, where $i\neq j$. Now we have two cases: when $f$ and $g$ are binomials associated with inner intervals that are disjoint or they share a vertex. For the first, we have $\gcd(\mathop{\rm in}\nolimits(f),\mathop{\rm in}\nolimits(g)=1$ trivially. For the second case, since the inner 2-minors $f$ and $g$ are of the form as shown in Figure 4 and $S(f,g)$ reduces to 00, so by Lemma 2.1 we obtain $\gcd(\mathop{\rm in}\nolimits(f),\mathop{\rm in}\nolimits(g))=1$ as well. So, for all $i,j\in[n]$ with $i\neq j$ and for all $f,g$ inner 2-minors of $I_{\mathcal{P}}$ such that $f$ is related to an inner interval of $\mathcal{P}_{i}$ and $g$ is related to an inner interval of $\mathcal{P}_{j}$, we have $\gcd(\mathop{\rm in}\nolimits(f),\mathop{\rm in}\nolimits(g))=1$. As a consequence, if $v\in V(\mathcal{P}_{i})\cap V(\mathcal{P}_{j})$ for some $i,j\in[n]$ with $i\neq j$, and $x_{v}\in\mathrm{supp}(\mathop{\rm in}\nolimits(f))$, for some inner 2-minor $f$ related to an inner interval of $\mathcal{P}_{i}$, then $x_{v}\notin\mathrm{supp}(\mathop{\rm in}\nolimits(g))$, for all inner 2-minors $g$ related to the inner intervals of $\mathcal{P}_{j}$. Let $A_{\mathcal{P}}=\{v\in V(\mathcal{P})\mid v\in V(\mathcal{P}_{i})\cap V(\mathcal{P}_{j}),\text{ for some }i,j\in[n],i\neq j\}$. Let $C_{\mathcal{P}}$ be the set of vertices $v\in A_{\mathcal{P}}$ such that $x_{v}\notin\mathrm{supp}(\mathop{\rm in}\nolimits(f))$, for all inner 2-minor $f$ related to an inner interval of $\mathcal{P}$. For each $i\in[n]$, define $D_{\mathcal{P}_{i}}$ the set of vertices $v\in V(\mathcal{P}_{i})\cap A_{\mathcal{P}}$ such that $x_{v}\notin\mathrm{supp}(\mathop{\rm in}\nolimits(f)),\text{ for all inner 2-minor }f\text{ related to an inner interval of }\mathcal{P}_{i}$. Denote $d_{i}=|D_{\mathcal{P}_{i}}|$. Set, for all $i\in[n]$, $S^{\prime}_{\mathcal{P}_{i}}=K[x_{a}:a\in V(\mathcal{P}_{i})\setminus D_{\mathcal{P}_{i}}]$. Then we have:
@@
Observe that for all $i\in[n]$ we have $S_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})=S^{\prime}_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})\otimes_{K}K[x_{v}\mid v\in D_{\mathcal{P}_{i}}]$. In particular, by [37, Lemma 5.1.11], $\mathrm{HP}_{S^{\prime}_{\mathcal{P}_{i}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}_{i}})}(t)=(1-t)^{d_{i}}\mathrm{HP}_{K[\mathcal{P}_{i}]}(t)$, and since $\mathcal{P}_{i}$ is a simple polyomino, by Proposition 1.3 we have $\mathrm{HP}_{S^{\prime}_{\mathcal{P}_{i}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}_{i}})}(t)=\frac{h_{\mathcal{P}_{i}}(t)}{(1-t)^{|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-d_{i}}}$. Note that $\sum_{i=1}^{n}|V(\mathcal{P}_{i})|=|V(\mathcal{P})|+|A_{\mathcal{P}}|$ and $\sum_{i=1}^{n}|\mathcal{P}_{i}|=|\mathcal{P}|$. Therefore, by [37, Lemma 5.1.11], we have:
@@
By the previous result, we argue that, if $\mathcal{P}$ is a simple collection of cells, the study of the $h$-polynomial of $K[\mathcal{P}]$ in terms of the switching rook polynomial of $\mathcal{P}$ can be reduced to investigate the Hilbert-Poincaré series of its connected components. In particular, if for all simple polyominoes [30, Conjecture 3.2] is true, then the same holds for all simple collections of cells.
@@
Let $\mathcal{P}$ be a simple collection of cells where every connected component $\mathcal{P}_{i}$ is a parallelogram polyomino. Then, by [30, Theorem 3.5], $h_{K[\mathcal{P}_{i}]}(t)$ is the switching rook polynomial of $\mathcal{P}_{i}$ for all $i\in[n]$, so $h_{K[\mathcal{P}]}$ is the switching rook polynomial of $\mathcal{P}$ by Corollary 2.3.
As a consequence, we generalize [32, Theorem 3.13 and Theorem 4.2] to simple thin collections of cells (see Figure 1c and 1d for an example). This is useful in the next sections.
@@
As a consequence of Corollary 2.3 and [32, Theorem 3.13], we get that $h_{K[\mathcal{P}]}$ is the rook polynomial of $\mathcal{P}$. Moreover, $\operatorname{reg}(K[\mathcal{P}])=r(\mathcal{P})$ since $K[\mathcal{P}]$ is a Cohen-Macaulay ring. Now we prove the result about Gorensteinness. Let $h_{K[\mathcal{P}]}(t)=r_{\mathcal{P}}(t)=\sum_{k=0}^{s}r_{k}t^{k}$, where $s=r(\mathcal{P})$. Since $K[\mathcal{P}]$ is a Cohen-Macaulay domain, it is known from [35] that $K[\mathcal{P}]$ is Gorenstein if and only if $r_{i}=r_{s-i}$ for all $i=0,\dots,s$, that is, $h_{K[\mathcal{P}]}(t)$ is palindromic. Assume that $\mathcal{P}$ consists of $n$ connected components and denote by $h_{K[\mathcal{P}_{i}]}(t)$ the $h$-polynomial of $K[\mathcal{P}_{i}]$ for all $i\in[n]$. Suppose that $\mathcal{P}$ has the $S$-property and we prove that $K[\mathcal{P}]$ is Gorenstein. Since $\mathcal{P}$ has the $S$-property, then each connected component $\mathcal{P}_{i}$ has the $S$-property as well. By [32, Theorem 4.2], $h_{K[\mathcal{P}_{i}]}(t)$ is palindromic for all $i\in[n]$, so $h_{K[\mathcal{P}]}(t)$ is palindromic as well, since $h_{K[\mathcal{P}]}(t)=\prod_{i=1}^{n}h_{K[\mathcal{P}_{i}]}(t)$ and the product of palindromic polynomials is palindromic. Hence it follows by [35] that $K[\mathcal{P}]$ is Gorenstein. Now, suppose that $K[\mathcal{P}]$ is Gorenstein and we show that $\mathcal{P}$ has the $S$-property. In particular the rook polynomial of $\mathcal{P}$, that is $h_{K[\mathcal{P}]}$, is palindromic and observe also that $r_{0}=1$ and $r_{1}=|\mathcal{P}|$. We assume by contradiction that $\mathcal{P}$ does not have the $S$-property. For all $i\in[n]$, let $h_{K[\mathcal{P}_{i}]}(t)=\sum_{j=0}^{s_{i}}r_{j}^{(i)}t^{j}$ be the $h$-polynomial of $K[\mathcal{P}_{i}]$, where $s_{i}=r(\mathcal{P}_{i})$ and observe that $r_{1}^{(i)}=|\mathcal{P}_{i}|$. Looking at the proof of case (b)$\Rightarrow$(c) of [32, Theorem 4.2], for all $k\in[n]$ such that $\mathcal{P}_{k}$ has not the $S$-property we have $r_{s_{k}}>1$ or $r_{s_{k}-1}>|\mathcal{P}_{k}|$. Therefore, it easily follows that $r_{s}>1$ or $r_{s-1}>\sum_{i=1}^{n}|\mathcal{P}_{i}|=|\mathcal{P}|$, that is a contradiction, since $h_{K[\mathcal{P}]}(t)$ is palindromic. ∎